ONE LEG IS 5 THE HYPOTENUSE IS 7 AND THE OTHER LEG IS 2X-2 WHAT ARE THE POSSIBLE VALUES OF X
Answer by rfer(16322)

49-25=24
sqrt 24=4.9
2x-2=4.9
2x=6.9
x=3.45
--------------
2*3.45-2=4.9